Grade 12th passAlgebra

Find:
15% of 250
Find:
15% of 250
Find:
15% of 250
Find:
15% of 250

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have,
= (15/100) × 250
= (15/10) × 25
= (15/2) × 5
= (75/2)
= 37.5

Thanks